OpenStudy (anonymous):

In the reaction, CH4 + 2O2 arrow CO2 + 2H2O, 890 kilojoules of heat are released. Using Einstein's equation, calculate the mass amount to which this energy equates. (1 J = 1 kg m2 s-2) and the speed of light = 3.0 x 108 m/sec (Points : 3) 9.9 x 1012 kg 9.9 x 10-12 kg 9.9 x 10-14 kg 1.1 x 10-6 kg

OpenStudy (abb0t):

Einsteins equation relating mass and energy is defined as: \(E = mc^2\) where: E = energy m = mass c = the speed of light: \(3.0 \times 10^8 \frac{ m }{ \sec }\) Use algebra to rearrange the formula to find mass. This means that energy is divided by the speed of light squared.